OpenStudy (anonymous):

when 5 is subtracted from a number the result is 16 ?

OpenStudy (anonymous):

21

OpenStudy (anonymous):

the answer is 21 do opposite 5 + 16 = 21

OpenStudy (anonymous):

hope that helped

OpenStudy (anonymous):

21 - 5 = 21

OpenStudy (anonymous):

x-5=16 add five on both sides x-5+5=16+5 x=21